What is logx+log(x-1)=log2? Solve for x
Answer by stanbon(75887)

logx+log(x-1)=log2
------------------------
log[x(x-1)] = log(2)
----
x(x-1) = 2
x^2-x-2 = 0
----
Factor:
(x-2)(x+1) = 0
Positive solution:
x = 2
=================
Cheers,
Stan H.